a room twice as long as the width. If the area of the room is 72 square feet, what is the perimeter.
Answer by amarjeeth123(569)

Let the width be x.
Then the length is 2x.
Area=length*width
2x*x=72
2x^2=72
x^2=36
x=6 feet.
The width of the room is 6 feet.
The length of the room is 12 feet.
The perimeter=2(length+width)=2(12+6)=2(18)=36 feet
The perimeter is 36 feet.